Could you help me to solve it, I tried many times and I got some answers but the website that Iam working on told me that " answer is wrong".
Find the slope, x-intercept, and y-intercept for the line 8x−3y−17=0 .
Note: Your answers must be decimals.
Thank You.

Answer by lwsshak3(11628)

Find the slope, x-intercept, and y-intercept for the line 8x−3y−17=0 .
Note: Your answers must be decimals.
***
8x−3y−17=0
This is an equation of a straight line of the standard form:y=mx+b, m=slope, b= y-intercept.
For given equation:
3y=8x-17
y=(8/3)x-17/3
..
slope=8/3=2.6667
y-intercept=-17/3=-5.6667
..
x-intercept
set y=0
8x-17=0
8x=17
x=17/8=2.1250
x-intercept=2.1250